OpenStudy (anonymous):

350/21 = 525/x I used to cross-multiply and then get x alone but this wasn't producing the right answer sometimes and gives me giant numbers to work with. Are there any easier methods to solve for x?

OpenStudy (anonymous):

if your looking for exact answers, then no, there arent really easier methods than cross multiplying.

OpenStudy (anonymous):

Ah I guess I probably just messed up some numbers when I got wrong answers. Thanks.

OpenStudy (anonymous):

525/(350/21) = 31.5, i think 31.5 is not a giant number

OpenStudy (anonymous):

I'll show you the steps I take. \[\frac{ 350 }{ 21} = \frac{ 525 }{ x }\]Cross multiply\[350x = 11025\]Get x alone by dividing 350 on both sides to get 31.5. The answer however needs to be a fraction and I got a decimal. How do I turn this decimal into a fraction?

OpenStudy (anonymous):

31.5 = 315/10

OpenStudy (anonymous):

x = 11025/350 = 315/10

OpenStudy (anonymous):

and then...

OpenStudy (anonymous):

simplify a bit more @suzi20

OpenStudy (anonymous):

i think he doesnt need mix fraction

OpenStudy (anonymous):

numerator and denominator are both divisible by 5... simplify fraction

OpenStudy (anonymous):

ok but he has to do it alone, we just give a little light in the darkness

OpenStudy (anonymous):

Absolutely! I love your approach.